(Intermediate Algebra: Rational Expressions)
4. For all r≠(+/-) 2, (r^2-5r+6)/(r^2-4) = ?
A. r-3/r+2
B. r-2/r+2
C. r-2/r+3
D. r+3/r-2
E. r+3/r+2
I'm confused, showing work would be great. I know its (A) but I don't know why.

Answer by Nate(3500)

(r^2-5r+6)/(r^2-4)
((r-3)(r-2))/((r-2)(r+2)) first, you simplify
Now, think about this. 1/1=1 and 2/2=1 and x/x=1 and (x-1)/(x-1)=1 ....
((r-3)(r-2))/((r-2)(r+2))
1(r-3)/(r+2)
so .... the answer is (A)